# Video2BEV: Transforming Drone Videos to BEVs for Video-based Geo-localization

### v1

@@ Abstract
Existing approaches to drone visual geo-localization predominantly adopt the image-based setting, where a single drone-view snapshot is matched with images from other platforms. Such task formulation, however, underutilizes the inherent video output of the drone and is sensitive to occlusions and environmental constraints. To address these limitations, we formulate a new video-based drone geo-localization task and propose the Video2BEV paradigm. This paradigm transforms the video into a Bird’s Eye View (BEV), simplifying the subsequent matching process. In particular, we employ Gaussian Splatting to reconstruct a 3D scene and obtain the BEV projection. Different from the existing transform methods, e.g., polar transform, our BEVs preserve more fine-grained details without significant distortion. To further improve model scalability toward diverse BEVs and satellite figures, our Video2BEV paradigm also incorporates a diffusion-based module for generating hard negative samples, which facilitates discriminative feature learning. To validate our approach, we introduce UniV, a new video-based geo-localization dataset that extends the image-based University-1652 dataset. UniV features flight paths at $30^{\circ}$ and $45^{\circ}$ elevation angles with increased frame rates of up to 10 frames per second (FPS). Extensive experiments on the UniV dataset show that our Video2BEV paradigm achieves competitive recall rates and outperforms conventional video-based methods. Compared to other methods, our proposed approach exhibits robustness at lower elevations with more occlusions.
11footnotetext: Corresponding author.
Figure 1: Typical failure cases for image-based drone geo-localization. For image queries, the core areas in ground-truth are occluded by another building, largely compromising the spatial matching. In contrast, video queries usually contain unoccluded frames in a circling flight, and thus could reflect a more comprehensive view of the target location. Drone video inputs are relatively robust to occlusions.
@@ 1 Introduction
Given the visual information captured by the drone, drone visual geo-localization aims to retrieve the image of the same location from another platform, e.g., satellite, which is typically associated with the off-line GPS metadata [54]. This capability enables drones to locate themselves even in GPS-unavailable areas, such as those between tall buildings or in rural regions. The common setting is based on the snapshot image from the drone as a query to search the matched location in the satellite-view candidate pool. Despite its potential, image-based drone geo-localization faces three primary challenges: (1) The single image from the drone view is often affected by occlusion and other environmental constraints. (2) The geometric transformation based on a single image usually introduces edge distortions and blurring. (3) Many locations share similar patterns, such as architectural styles, making it challenging to distinguish between them using a single image alone.
This paper attempts to address three challenges in drone visual geo-localization. First, existing approaches [54, 43, 42, 5] predominantly adopt the image-based setting, where a single drone-view snapshot is matched with images from other platforms. As shown in Fig. 1, such task formulation underutilizes the inherent video output of the drone and is sensitive to occlusions and environmental constraints. To mitigate this, we formulate a new video-based drone geo-localization task, which harnesses the drone video as input to capture the target location comprehensively. By controlling the drone to achieve unobstructed viewpoints, the video input significantly reduces the impact of occasional occlusions present in the single frame, thereby improving the performance on all our re-implemented methods (see  Fig. 2).
Figure 2: Performance comparisons of leveraging image data (Image2Image) or video data (Video2Image or Image2Video) with methods including LPN [42], FSRA [3], DWDR [43], and Sample4Geo [5]. We report the Average Precision (AP) metric. For a fair comparison, we keep the same number of data in the gallery. We could observe that all our re-implemented methods arrive at better performance when adopting video query or gallery.
Second, geometric transformations based on a single image often introduce edge distortion and blurring (see Fig. 3). Existing approaches [30, 45] usually apply the transformation pre-processing to align the input images collected from different platforms. Such transformations coarsely align the images, thereby easing spatial comparison for deeply-learned models. For instance, Shi et al. [30] propose transforming satellite views to ground views via the polar transformation, which generally aligns the layout of the transformed images with fisheye ground-view photos. Similarly, Wang et al. [45] manipulate panoramic ground views to satellite views via the spherical transformation, simplifying the comparison with satellite-view data. However, both the pre-processing transformation methods suffer from spatial distortion and blurring, leading to incomplete image-level alignment. Since the input of our task is the video with different viewpoints along the flight, we propose a Video2BEV paradigm that transforms the video into a Bird’s Eye View (BEV). Specifically, we employ Gaussian Splatting [12] to reconstruct a 3D scene and obtain the BEV projection. Unlike existing transformation methods, such as polar and spherical transformations, our transformation is not based on a single 2D image but a 3D scene derived from videos. Therefore, the core area of our BEVs preserve more fine-grained textures without significant distortion or blurring.
Third, many locations share similar patterns, such as architectural styles, making it challenging to distinguish between them using a single image alone. To address this, we resort to the drone video input in this work and conduct more discriminative feature learning. To enhance model scalability and discriminative capability, our Video2BEV paradigm incorporates a diffusion-based module for generating hard negative samples. This module generates appearance-similar BEVs, motivating the model to focus on fine-grained details to discriminate diffusion-generated BEV hard negatives and BEVs derived from videos.
@@
In an attempt to overcome the aforementioned challenges, this paper (1) adopts the video setting, (2) transforms video to BEVs derived from reconstructed 3D scenes, (3) introduces a diffusion model to generate hard negatives. Given the lack of a video-based drone geo-localization benchmark, we introduce UniV, a new video-based geo-localization dataset that extends the image-based University-1652 dataset. UniV features flight paths at two different elevation angles with increased frame rates up to 10 frames per second (FPS). Extensive experiments on the UniV dataset show that our Video2BEV paradigm achieves competitive recall rates and outperforms conventional video-based methods. Compared to other methods, our proposed approach exhibits greater robustness at lower elevations with more occlusions. To summarize, the key contributions of our work are as follows:
  • We present a Video2BEV paradigm that transforms drone-view videos to Bird’s Eye Views (BEVs), easing subsequent matching with satellite images. In particular, our Video2BEV introduces 3D Gaussian Splatting for geometric projection, while also including a hard negative generation module to learn from diverse BEVs and satellite images.
  • Given the lack of a video-based drone geo-localization benchmark, we introduce a new video-based geolocalization dataset, UniV, with two elevation angles and 10 FPS. This dataset contains drone videos, satellite images, and ground images, and is closer to real-world scenarios containing typical cases, e.g., occlusions.
  • Extensive experiments on the proposed UniV dataset show: (1) Image-based queries are highly sensitive to occlusions and environmental changes, such as occlusions, whereas video-based queries exhibit greater robustness (see Fig. 2). (2) The proposed Video2BEV, which incorporates geometric transformations, achieves 96.80 AP on Drone Video $\rightarrow$ Satellite, significantly outperforming other existing methods. We observe a similar result on the drone video with the lower elevation angle, which exposes more occlusions.
@@
Table 1: (a) Dataset comparisons between UniV and other visual geo-localization datasets. G, S, and D denote ground-view, satellite-view and drone-view, respectively. We enable video modality and add another common elevation angle of drone flight. (b) Elevation angles $\theta$ illustration. Top panel shows $\theta$ = $45^{\circ}$ and bottom panel displays $\theta$ = $30^{\circ}$. With a lower elevation angle, the new flight captures the target building with wider Field of View (FoV) but more occlusions, thereby posing more challenges for drone visual geo-localization.
Datasets Platforms #data/location Modality Elevation
@@
GeoText-652 G, S, D (16 + 1 + 54) images + Image + $45^{\circ}$
180 texts Text
UniV G, S, D (16 +1) image + Video $30^{\circ}$, $45^{\circ}$
2 videos
@@ 2 Related Work
Visual Geo-localization. Visual geo-localization is to locate the position via the visual information, which is usually formulated as a sub-task of image retrieval [54]. The ancient people could check the surroundings to find their location via the paper map. Similarly, visual geo-localization searches the relevant position candidates recording in other platforms, e.g., satellite. The primary challenge of this task is due to the inherent difference among multiple platforms, i.e., appearance changes due to various viewpoints [30]. Therefore, previous methods focus on alignment to establish spatial correspondence between these platforms, which can be coarsely divided into two families, i.e., image-level alignment and feature-level alignment. For image-level alignment, Shi et al. [30] first propose to leverage polar transformation to warp satellite images to the ground view. Other methods [45, 13] transform panoramic ground images to the satellite view by proposed spherical transformation. Regmi et al. [26] utilize generative adversarial networks to synthesize aerial images for matching with ground images. However, the transformed images often suffer from severe distortion, blurring, and unrealistic content, which hinders fine-grained alignment between the two distinctive viewpoints. For implicit feature-level alignment, prior works [54, 5, 57, 36, 50, 51, 21] tend to rely on the power of neural networks or loss functions to align features between the query and gallery. Other methods focus on the explicit alignment of features from different views. Among these, some methods [18, 31, 32] encode orientation information and utilize it for feature-level alignment. Some methods propose to establish key-point [16, 35] or region alignment [3, 49, 48, 46] in the feature level. Previous works overlook the viewpoint variation within drone-view videos, treating them in an image-based setting. In this work, we aim to unlock the potential of videos captured by a drone and mitigate view disparity via the transformation without obvious distortion or blurring.
Video Understanding. The early works, e.g., Carreira et al. [2] and Shen et al. [29], leverage the two-stream 3D convolution network that combines video data with corresponding optical flow. In this way, motion information is adopted as the guidance for the fusion of different frames. Instead of adopting the 3D convolution, Feichtenhofer et al. [7] propose to deploy a two-branch 2D convolution network to fuse spatial semantic information and motion information together. Recently, attention mechanisms have shown their priority across many tasks [1, 40, 19, 4]. However, the computational cost of processing video data is substantial [8]. Many methods focus on designing efficient attention operations and training strategies. For efficient attention operations, Liu et al. [20] propose a 3D shifted window mechanism based on multi-head attention to fuse inherent complementary information of videos. Son et al. [34] propose CNN-based attention, modeling the ordering reliance in video frames and capturing the long-term dependency. For effective training strategies, Tong et al. [38] introduce mask image modeling [9] into video data in a self-supervised manner, leveraging complementary information across different frames. To keep task-specific and shared knowledge across different tasks, Peirone et al. [24] treat video understanding as a unified task and propose to train task-specific heads and a cross-task backbone simultaneously. Different from the daily videos, drone videos for geo-localization typically contain multi-view information for the target location [44, 22]. Therefore, rather than adopting the off-the-shelf video backbone, we propose a Video2BEV transformation to leverage the 3D geometric correspondences and enable a straightforward spatial alignment for matching.
@@ 3 The UniV Dataset
Given the lack of a video-based drone geo-localization benchmark, we collect a new dataset dubbed UniV involving the video modality. We follow the building information and the protocol of the existing University-1652 dataset [54]. The UniV dataset encompasses 1,652 buildings in 72 universities from three platforms, i.e., ground, satellite, and drone cameras. In particular, the UniV dataset contains 16 ground-view images and 1 satellite-view image for each building and the training set of UniV dataset contains 701 buildings, while the test set in the UniV dataset includes other 951 buildings. There are no overlapping buildings between the training and test sets. The proposed UniV dataset is different from the image-based University-1652 and other datasets in two primary aspects, i.e., modality and elevation-angle expansions (see  Tab. 1).
@@
Modality Expansion. Existing datasets [47, 41, 18, 56] collect data from two platforms, e.g., satellite and ground cameras. Although some datasets [54, 17, 37] include drone or aerial views, they still collect data in the image format. We adopt similar operations as the University-1652 dataset but collect drone-view data in video format. Specifically, we leverage the 3D engine of google earth [23] to simulate the movement of a drone-view camera in the real world. To collect video data containing both scale and viewing-point variations, we leverage the dynamic viewpoints within the 3D engine and set the moving viewpoints along a spiral curve for moving around the target building three circles, closely approximating real-world drone flights. Video data are collected in 30 frames per second. Considering the video redundancy, in practice, we subsample videos along the temporal dimension, resulting in frame rates of 2, 5, and 10 for further processing.
@@
Elevation-angle Expansion. Conventional datasets [54, 17, 37] collect drone or aerial data in a fixed elevation angle, i.e., $45^{\circ}$, which does not fully simulate the real-world use cases. Therefore, we add one new synthetic flying path at another common setting, i.e., a lower elevation angle $30^{\circ}$. The new flying path poses two new challenges (see Tab. 1). First, drones flying at a $30^{\circ}$ elevation angle capture scenes that include the target building and more surrounding areas, providing a wider Field of View (FoV). At the same time, it introduces disruptions for the center target building during matching. Second, there are more occluded cases, which lay over the core areas of the target building. It poses challenging to mine the discriminative frames in the video, while it is easier in the same location captured at a $45^{\circ}$ elevation. Therefore, the proposed dataset could further evaluate the robustness of the method against more disruptions and occlusions, which is closer to real-world drone usage.
@@
Discussion. The contribution to the community. The key difference between the UniV dataset and existing datasets [54, 17, 37] lies in the modality expansion of videos (see Tab. 1), facilitating the development of robust drone visual geo-localization. A single image provides limited information about the corresponding building. When core areas of the buildings are occluded, single-image queries can not produce reliable matching results (see Fig. 1). In such cases, the video contains both occluded and unoccluded frames. One frame may contain core-area information to complement another frame and together they can provide robust and complete information required for drone visual geo-location. In this way, all re-implemented methods perform better when adopting video data (see Fig. 2). Moreover, the UniV dataset also introduces a new real-world challenge for drone visual geo-localization. The new elevation angle of $30^{\circ}$ is typical in real-world flights***The United States and the United Kingdom allow drone flights up to 400 feet; China restricts drones up to 120 meters.. The $30^{\circ}$ elevation angle faces more occlusion cases (see Tab. 1), simulating outputs of real-world drone flights.

### v3

@@ Abstract
Existing approaches to drone visual geo-localization predominantly adopt the image-based setting, where a single drone-view snapshot is matched with images from other platforms. Such task formulation, however, underutilizes the inherent video output of the drone and is sensitive to occlusions and viewpoint disparity. To address these limitations, we formulate a new video-based drone geo-localization task and propose the Video2BEV paradigm. This paradigm transforms the video into a Bird’s Eye View (BEV), simplifying the subsequent inter-platform matching process. In particular, we employ Gaussian Splatting to reconstruct a 3D scene and obtain the BEV projection. Different from the existing transform methods, e.g., polar transform, our BEVs preserve more fine-grained details without significant distortion. To facilitate the discriminative intra-platform representation learning, our Video2BEV paradigm also incorporates a diffusion-based module for generating hard negative samples. To validate our approach, we introduce UniV, a new video-based geo-localization dataset that extends the image-based University-1652 dataset. UniV features flight paths at $30^{\circ}$ and $45^{\circ}$ elevation angles with increased frame rates of up to 10 frames per second (FPS). Extensive experiments on the UniV dataset show that our Video2BEV paradigm achieves competitive recall rates and outperforms conventional video-based methods. Compared to other competitive methods, our proposed approach exhibits robustness at lower elevations with more occlusions.
[…]
Figure 1: Typical failure cases for image-based drone geo-localization. For image queries (left), the core areas in ground-truth are occluded by another building, largely compromising the spatial matching. In contrast, video queries (right) usually contain unoccluded frames in a circling flight, and thus could reflect a more comprehensive view of the target location.
@@ 1 Introduction
Drone visual geo-localization aims to retrieve images of the same location from another platform, such as satellite, using visual information captured by the drone. This process is typically supported by off-line GNSS metadata [74], enabling drones to self-localize even in GNSS-denied environments, such as urban canyons or rural areas. The prevailing approach follows an image-based matching paradigm [74, 59, 58, 8, 6, 31], where a single drone-captured snapshot serves as the query to retrieve the corresponding location from the satellite-viewed candidate pool. However, despite the advancements in image-based paradigms, two critical limitations persist. Both are due to drone flight height regulations [52, 2, 19]. (1) Drones typically operate at lower altitudes in cluttered environments, resulting in significant occlusions in the captured images from buildings, trees, and other foreground objects. Such occlusion can lead to a substantial loss or degradation of visual information in the drone image captured from a single viewpoint, making it difficult to establish accurate correspondences with satellite imagery. As shown in Fig. 1 (left), the core areas in the query are entirely obstructed by surrounding buildings. (2) Similarly, due to height limitation, drones typically capture images at oblique angles, while satellite images are predominantly acquired from a top-down perspective. The significant viewpoint disparity between the drone and satellite perspectives further increases the difficulty of matching.
@@
To address the limitations of the image-based geo-localization paradigm, we formulate a new video-based drone geo-localization task and propose a corresponding paradigm named Video2BEV, which leverages drone videos and transforms them into Bird’s-Eye View (BEV) representations for drone-satellite matching. (1) Different from traditional single-view image approaches, our Video2BEV paradigm resorts to the multi-view nature of video to recover occluded regions and improve matching robustness. As shown in Fig. 1 (right), even though core areas of the target location are occluded in a certain drone-captured frame, we still can recover such core areas from other frames with different viewpoints. (2) In this work, we reconstruct the view in the BEV format, since BEV representation aligns with the satellite’s top-down viewpoint, thus reducing the inter-platform discrepancy. To convert input image into a calibrated format, existing methods usually apply 2D geometric transformations [41, 61], but suffer from spatial distortion and blurring (see Fig. 2 (a, b)). Inspired by the success of 3D Gaussian Splatting (3DGS) in reconstruction [4], we introduce a new 3D-aware transformation. In particular, we leverage 3DGS to reconstruct the 3D scene based on the multi-view snapshots from the drone video, and then obtain the BEV representation via projection. As shown in Fig. 2 (c), the BEV generated by our Video2BEV transformation exhibits fine-grained textures with minimal distortion or blurring, thus facilitating the subsequent inter-platform matching. Furthermore, considering the nearby location with a similar visual appearance, the proposed Video2BEV paradigm further incorporates a diffusion-based hard negative synthesis module. This module generates BEV representations that retain original semantic content but with different fine-grained discrepancies, serving as hard negatives during training. By incorporating these challenging samples, the model learns to discriminate intra-platform samples from highly similar yet geographically distinct locations.
Finally, to support the video-based geo-localization task, we introduce a new dataset UniV with 2 drone videos per location accompanying with 16 ground-view snapshots and 1 satellite-view image, which is closer to the real-world deployment. With the help of unobstructed frames, the video input significantly reduces the impact of occasional occlusions present in the single image, thereby improving overall performance on all our re-implemented methods. For instance, LPN [58] receives +24.46% AP increment (see  Fig. 3). In brief, our contributions are:
Figure 3: Performance comparisons of leveraging image data (Image2Image) or video data (Video2Image or Image2Video) with different methods including LPN [58], FSRA [6], DWDR [59], and Sample4Geo [8]. We report the Average Precision (AP) metric. For a fair comparison, we keep the same number of data in the gallery. We could observe that all re-implemented methods achieve better performance when adopting video query or gallery.
  • We formulate a new video-based geo-localization task and propose the Video2BEV paradigm that transforms drone-view videos into BEV representations with the assistance of 3DGS, simplifying the subsequent inter-platform matching process. To further enhance the intra-platform representation learning, we introduce a diffusion-based hard negative sample synthesis module, which generates challenging training samples to expand data diversity and improve discriminative capability.
  • To validate the video-based drone geo-localization task, we introduce a new dataset, called UniV, with 3,304 drone flight videos with corresponding satellite and ground-view images. Our experiments reveals two insights: (1) Video-based data shows consistent performance advantages over single-frame image retrieval across multiple metrics (see Fig. 3). (2) The proposed Video2BEV achieves 96.80 AP on Drone Video $\rightarrow$ Satellite, outperforming other competitive methods. Furthermore, the trained model shows strong generalization capabilities, maintaining 91.50 AP when tested on the unseen real-world dataset, i.e., SUES-200, without fine-tuning.
@@
Table 1: (a) Dataset comparisons between UniV and other visual geo-localization datasets. G, S, and D denote ground-view, satellite-view, and drone-view, respectively. We enable video modality and add another common elevation angle of drone flight. (b) Elevation angles $\theta$ illustration. Top panel shows $\theta$ = $45^{\circ}$ and bottom panel displays $\theta$ = $30^{\circ}$. With a lower elevation angle, the new flight captures the target location with wider Field of View (FoV) but more occlusions, thereby posing more challenges for drone visual geo-localization.
Datasets Platforms #data per location Modality Elevation
@@
SUES-200 [77] S, D (1 + 50 $\times$ 4) images Image $45^{\circ}\sim 70^{\circ}$
@@
GeoText-1652 [5] G, S, D (16 + 1 + 54) images + 180 texts Image + Text $45^{\circ}$
UniV G, S, D (16 +1) images + 2 videos Image +Video $30^{\circ}$, $45^{\circ}$
@@ 2 Related Work
Image-based Geo-localization. Image-based geo-localization, which is usually regarded as a sub-task of image retrieval, applies image query to determine locations [63, 71]. The primary challenge of this task is the large appearance discrepancy due to different viewpoints across platforms, including ground [17], satellite [11], and drone [9, 76]. Previous methods can be coarsely divided into two families: image-level alignment and feature-level alignment. (1) For image-level alignment, Shi et al. [41] leverage polar transformation to warp satellite images to the ground view. Similarly, Wang et al. [61, 67] transform ground images to the satellite view. Regmi et al. [36] synthesize aerial images from ground images to facilitate matching with satellite view via Generative Adversarial Networks (GANs). Tian et al. [49] employ GANs to transform drone-view images into satellite-view images. Andrea et al. [54] convert drone-view images to ground views through 3D reconstruction but the output is with distortions. (2) For feature-level alignment, Dai et al. [6] and Wang et al. [58] establish feature alignment in a region-correspondence manner. Some methods [26, 46] focus on key-point alignment. Other methods focus on improving the discriminative ability of neural networks with tailored modules, such as lite-transformer encoder [62], layer-to-layer attention block [79, 65], adaptive integration module [47], strong backbones [66], adaptation information consistency module [23], spatially-adaptive denormalization [60] and well-designed loss functions, e.g., semantic augmentation loss [70], contrastive loss [71, 31, 22, 8], dynamic weighted decorrelation regularization [59], peer learning [69], instance loss [74] and the optimal transport [43]. Additionally, other methods [37, 28, 42] fuse the extra orientation meta-information from GNSS with extracted features. However, previous methods overlook the viewpoint variation within drone-view videos, treating them in an image-based setting. Our method is among the early attempts to leverage the viewpoint variation of drone-view videos to transform drone-view video to Bird’s-Eye View (BEV), thereby reducing the viewpoint disparity between drone and satellite views.
Video-based Geo-localization While video understanding has been a focus of the computer vision community for decades, the problem remains challenging due to the complexity added by the time dimension and the volume of data. Early works [3, 40, 12] leverage two-stream convolution networks to fuse spatial semantic information with motion information. Subsequently, attention mechanisms [7, 13, 34] have been introduced for long-term video understanding, such as vanilla self-attention [1, 55], shift window [29, 30], masked auto-encoder [14, 51], and local spatiotemporal attention [45]. Recently, large language models [48] have also shown their superiority in video understanding. In visual geo-localization, videos contain more visual information captured through the camera’s trajectory, which can provide more comprehensive information compared to images. Vyas et al. [57] are the first to collect ground-view data in video format and propose a hierarchical approach for processing clips of ground-view videos. Regmi et al. [37] leverage the geo-temporal proximity between the ground-view videos and GNSS locations to extract coherent features from videos. Expanding to a global scale, Kulkarni et al. [21] introduce a large-scale ground-view video dataset for worldwide geo-localization. Different from the ground-view videos, drone-view videos typically contain multi-view and multi-scale information for the target location [32]. In this paper, we collect drone-view data in video format and propose a video-based geo-localization dataset. Rather than adopting the off-the-shelf video backbone, we propose a Video2BEV transformation to leverage the 3D geometric correspondences and enable a straightforward spatial alignment for matching.
@@ 3 The UniV Dataset
Given the lack of a video-based drone geo-localization benchmark, we collect a new dataset dubbed UniV involving the video modality. We follow the location information and the protocol of the existing University-1652 dataset [74]. The UniV dataset encompasses 1,652 locations in 72 universities from three platforms, i.e., ground, satellite, and drone cameras. In particular, the UniV dataset contains 16 ground-view images and 1 satellite-view image for each location and the training set of UniV dataset contains 701 locations, while the test set in the UniV dataset includes other 951 locations. There are no overlapping locations between the training and test sets. The proposed UniV dataset is different from the image-based University-1652 and other datasets in two primary aspects, i.e., modality and elevation-angle expansions (see  Tab. 1).
@@
Modality Expansion. Existing datasets [64, 56, 28, 78] collect data from two platforms, e.g., satellite and ground. Although some datasets [74, 50, 77] include drone views, collected data is still in an image format. We adopt similar operations as the University-1652 but collect drone-view data in a video format. Specifically, we leverage the 3D engine of google earth [33] to simulate the real-world movement of a drone equipped with a camera. To collect video data containing various scales and viewpoints, we leverage the dynamic viewpoints within the 3D engine and set the moving viewpoints along a spiral curve for moving around the target location in three circles, closely approximating real-world drone flights. All videos are collected in 30 frame rate. Considering the video redundancy, in practice, we subsample videos along the temporal dimension, resulting in frame rates of 2, 5, and 10 for further processing.
@@
Elevation-angle Expansion. Conventional datasets [74, 27, 50] collect drone data in a fixed elevation angle, i.e., $45^{\circ}$, which does not fully simulate the real-world cases. Therefore, we add one new synthetic flying path at another common setting, i.e., a lower elevation angle $30^{\circ}$. The new flying path poses two new challenges (see Tab. 1). First, drones flying at a $30^{\circ}$ elevation angle capture scenes that include the target location and more surrounding areas, providing a wider Field of View (FoV), thus introducing disruptions for the center target location during matching. Second, flight paths at a lower elevation angle lead to more occluded cases, which lay over the core areas of the target location. It poses challenges for mining the discriminative frames in the video, whereas it becomes easier when captured at a $45^{\circ}$ elevation. Therefore, the proposed dataset could further evaluate the robustness of methods against more disruptions and occlusions, which is closer to real-world drone geo-localization usage.
@@
Discussion. The contribution to the community. Different from existing datasets [74, 27, 50], the proposed UniV expands the modality from image to video (see Tab. 1), facilitating the development of robust drone visual geo-localization. A single image provides limited information about the corresponding location. When core areas of the location are occluded, single-image queries can not produce reliable matching results (see Fig. 1). In such cases, the video contains both occluded and unoccluded frames. One frame may contain core-area information to complement another frame and together they can provide robust and complete information required for drone visual geo-location. In this way, all re-implemented methods perform better when adopting video data (see Fig. 3). Moreover, the UniV dataset also introduces a new real-world challenge for drone visual geo-localization. The new elevation angle of $30^{\circ}$ is typical in real-world flightsThe United States and the United Kingdom allow drone flights up to 400 feet; China restricts drones up to 120 meters.. The $30^{\circ}$ elevation angle faces more occlusion cases (see Tab. 1), simulating outputs of real-world drone flights.